Captain John Stapp of the us air force. Tested the human limits of acceleration by riding on a rocket sled of his own design,known as the geewhiz. What netforce would be required to accelerate the 82-kg Stapp at 450m/s2(the highest accelerated tested by Stapp?)
1
Expert's answer
2020-12-03T10:25:16-0500

We can find the net force from the Newton's Second Law of Motion:


"F_{net}=(m_{sled}+m_{Stapp})a,""F_{net}=(82\ kg+82\ kg)\cdot 450\ \dfrac{m}{s^2}=73800\ N."

Answer:

"F_{net}=73800\ N."